OpenStudy (anonymous):

These triangles are similar. What is the value of x? A) 6 cm B) 7 cm C) 8 cm D) 9 cm

OpenStudy (anonymous):



OpenStudy (anonymous):

Set this up as a proportion \( \huge \frac{18}{10} =\frac{x}{5} \) Now just solve for x

OpenStudy (anonymous):

What is the value of x in the equation I gave you?

OpenStudy (anonymous):

In similar triangles, the ratios of corresponding sides are equal. Therefore, \[\frac{ x }{ 18 }=\frac{ 5 }{ 10 }\] Solve for x

OpenStudy (anonymous):

yeah breanna solve for x

OpenStudy (anonymous):

dont you cross multiply?? sorry im so dumb

OpenStudy (anonymous):

Yes cross multiply and then divide by 10

OpenStudy (anonymous):

is it 9?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

thanks you guyzzz =^w^=

OpenStudy (anonymous):

YW